# Systematic assessment of various universal machine-learning interatomic potentials

### v1

@@ Abstract
[Abstract]Machine-learning interatomic potentials have revolutionized materials modeling at the atomic scale. Thanks to these, it is now indeed possible to perform simulations of ab initio quality over very large time and length scales. More recently, various universal machine-learning models have been proposed as an out-of-box approach avoiding the need to train and validate specific potentials for each particular material of interest. In this paper, we review and evaluate five different universal machine-learning interatomic potentials (uMLIPs), all based on graph neural network architectures which have demonstrated transferability from one chemical system to another. The evaluation procedure relies on data both from a recent verification study of density-functional-theory implementations and from the Materials Project. Through this comprehensive evaluation, we aim to provide guidance to materials scientists in selecting suitable models for their specific research problems, offer recommendations for model selection and optimization, and stimulate discussion on potential areas for improvement in current machine-learning methodologies in materials science.
@@
Materials simulations at the atomic scale are the backbone of computational materials design and discovery. They rely on the Born Oppenheimer approximation, in which the electrons follow the nuclear motion adiabatically, so that the potential governing the nuclei consists of the electronic energies as a function of the nuclear positions, called "potential energy surface" (PES). Knowledge of the PES allows the identification of stable and metastable atomic configurations from minimum energy search, or the determination of materials properties as thermodynamical averages from molecular dynamics simulations 1. The utility of atom-based materials simulations is thus intimately related to the generation of accurate PESs, which has been possible in the last decades thanks to the advent of density-functional theory (DFT) 2, 3, 4, 5. Nonetheless, this ab initio approach relies on the quantum mechanical solution of the electronic problem whose computational cost scales cubically with system size and can therefore become unaffordable in various significant cases of technological interest such as amorphous solids, interfaces, surfaces, etc. At the other end of the simulation approaches, parametrized approximations of the Born-Oppenheimer PES, known as empirical analytical potentials, or force fields, or "classical” interatomic potentials, have been widely used especially for large-scale materials studies 6. Unfortunately, in particular when complex electron interactions are involved (as in chemical reactions or phase transitions) these approaches cannot usually achieve DFT accuracy, and in addition they have limited applicability and transferability. They cannot therefore be considered as a drop-in replacement for standard ab initio methods. In this context, machine-learning interatomic potentials (MLIPs) have emerged as an in-between solution with computational cost similar to the empirical analytical potentials, but with the promise of achieving an accuracy comparable to DFT hence enabling accurate simulations over very large time and length scales 7. The key difference with respect to the empirical potentials is that the interatomic potentials are now directly obtained via a highly non linear fit of a set of input/target data (in general, at the DFT accuracy), without any a priori assumption on their analytical form 8. In the original formulation, building a MLIP consists of generating a dataset of atomic configurations for the specific material under study, and training (and subsequently validating) the MLIP on these data based on the accurate prediction of some target metrics as, e.g., energies, forces, and stresses 9, 10, 11, 12, 8, 13, 14, 15, 16. This process is highly material-dependent, and usually requires a significant human and computational effort 17. Obtaining an accurate description of the PES without the need for costly DFT computations, but also covering all possible chemical and structural spaces, would be the holy grail of MLIPs.
Graph neural network based methods can be particularly useful for generalizability thanks to their property of “learning locally” that makes the resulting potential less material-dependent 8. Indeed, the first “universal” MLIP (uMLIP), the MEGNet 18 model, exploited a graph network architecture. It was trained on 60000 inorganic crystals in their minimum energy configurations from the Materials Project (MP) 19 database, which covers the majority of the elements of the periodic table (89) and is based on the Perdew-Burke-Enzerhof (PBE) exchange-correlation functional 20. This model could provide the formation energy as well as a number of other properties. In order to predict forces and stresses, various other models were subsequently developed relying on different datasets. In particular, the M3GNet 21 and CHGNet 22 models relying on equivariant graph neural network architectures were both developed using snapshots from DFT relaxations of the MP structures. The publication introducing CHGNet was also the opportunity for releasing the Materials Project Trajectory (MPtrj) dataset 22, including the DFT calculations for more than 1.5 million atomic configurations of inorganic structures. By including magnetic moments in the training properties, CHGNet aimed at a better description of chemical reactions as charged states influence how atoms connect with others through chemical bonds 22. A re-implementation of M3GNet, MatGL, has been built on the Deep Graph Library and on PyTorch in order to improve usability and scalability 23. The ALIGNN-FF model 24 was developed to model a diverse set of materials with any combination of 89 elements from the periodic table but relying on a different database of inorganic crystals, JARVIS-DFT 25, which is based on the optB88vdW exchange-correlation functional 26. While writing the present paper, the MACE-MP-0 27 relying on the MACE architecture 28 was also proposed. It was trained on the MPtrj dataset and showed outstanding performance on an extraordinary range of examples from quantum-chemistry and materials science 27. Two proprietary models relying on very large databases have also been developed: the GNoME 29 model exploiting the NequIP architecture 15 and the PFP model 30, 31 exploiting the TeaNet architecture 32. On the one hand, GNoME was trained on a database obtained from a complex active learning workflow of the original MP data, resulting in a number of inorganic structures 100 times larger than the MPtrj. On the other hand, the PFP was trained on a large dataset which initially included 10 DFT configurations covering 45 elements 30 and was further extended to cover 72 elements at the moment of the publication 31, with an expected rise to 94 including rare-earth elements and actinides. Finally, it is worth mentioning that uMLIPs have also been developed specifically for organic molecules 33, 34, 35, and for metal alloys 36.
@@
In this paper, we conduct a comprehensive review and evaluation of five different graph neural network (GNN)-based uMLIPs: M3GNet 21, CHGNet 22, MatGL 21, 23, MACE-MP-0 27 (called for simplicity MACE in what follows), and ALIGNN-FF 37 (called for simplicity ALIGNN in what follows) . They have demonstrated the possibility of universal interatomic potentials that may not require retraining for new applications. CHGNet is different from the other models considered here since it takes magnetic moments into account. The evaluation uses three different datasets: for the equation of state comparison, we use the set of theoretical structures employed in Ref. 38, for the phonon calculations we use the crystalline structures considered in Ref. 39 that have been relaxed with Abinit 40, 41 norm-conserving pseudopotentials 42, 43 and the PBEsol exchange-correlation functional while for all the other tests, we use the VASP-relaxed structures from the Materials Project (MP) 19, 44.
@@
The paper is organized as follows. In Sec 2 we test and discuss the quality and transferability of the chosen uMLIPs in different types of calculations, and using the three different datasets as explained above. The tests include the calculation of the equation of state (Sec. 2.1), the evaluation of formation energies and the optimization of structural parameters (Sec. 2.2), and the calculation of phonon bands (Sec. 2.3). The conclusions are provided in Sec. 3 and the details of the methods employed are given in Sec.4.
@@ 2.1 Equation of state and comparison with all-electron results
As a first test, we use the protocol for the equation of state (EOS) detailed in Ref. 38, where a high-quality reference dataset of EOS for 960 cubic crystal structures is generated by employing two all-electron (AE) codes. This dataset includes all elements from Z = 1 (hydrogen) to Z = 96 (curium). For each element, four mono-elemental cubic crystals (unaries structures) are considered, in the face-centered cubic (FCC), body-centered cubic (BCC), simple cubic (SC), and diamond crystal structure, respectively. Besides, six cubic oxides (oxides structures) are included for each element X, with chemical formula XO, XO, XO, XO, XO, and XO, respectively. While in Ref. 38 this dataset is used to gauge the precision and transferability of nine pseudopotential-based ab initio codes, we use it here to assess the five tested uMLIPs.
@@
In Fig. 1 the relative errors on the uMLIPs predictions for the equilibrium volume , the bulk modulus , and its derivative with respect to the pressure obtained based on a fit 38 of the EOS, are compared with the analogous errors from selected ab initio methods employed in Ref. 38. FLEUR 45 and WIEN2K 46 are AE codes while VASP implements the projector augmented-wave method (PAW) method 47 with a planewave basis set. The errors reported in Fig. 1 are calculated respect to the average of the AE methods considered here, i.e., FLEUR 45 and WIEN2K 46.
@@
Also, as in Ref. 38, the EOS computed with two different computational approaches and ( and ) are compared through two metrics and . The first is a renormalized dimensionless version of the metrics introduced in Ref. 48:
@@
where the index runs over the explicit calculations of for the different methods and is the integral average of over the considered volume range. The second metric depends directly on the physically measurable quantities , , and . It captures the relative deviation for each of these three parameters between the two computational approaches and :
@@
where , , and are appropriately chosen weights (see Ref. 38). Here, and are a uMLIP and the average of the AE methods considered here, i.e., FLEUR 45 and WIEN2K 46, respectively. Heatmaps of the periodic table with the values of the comparison metrics and obtained with all the different uMLIPs are reported in Fig. 2 in the Supplemental Material.
@@
It should be noted that most structures in the dataset used in this EOS test are not stable in nature. Therefore, it is very likely that these configurations were not included in the dataset used to train the uMLIPs. As a consequence, it is not surprising that the uMLIPs are not able to predict the correct energy versus volume curve for a significant fraction of systems (Figs. 1-5 in the Supplementary Material). Nevertheless, even for those systems for which a physical EOS is obtained, the precision and transferability is still far from the one that can be achieved with state–of-the-art pseudopotential-based ab initio techniques, as can be observed in Fig. 1. This is a very stringent test for uMLIPs, especially given the effort made in Ref. 38 to improve the precision and the transferability of pre-existent pseudopotential tables used for ab initio calculations. Yet these results suggest that uMLIPs predictions should be taken with some caution and, if possible, validated a posteriori via ab initio calculations, especially if the chemical/physical environment under study is not properly included in the training dataset. The precision of uMLIPs can be improved after retraining the model by including additional ab initio data capturing the chemical/physical configurations under investigation. However, despite their relevance, these topics are beyond the scope of the present work and are left for future investigation.
2.2 Atomic and lattice relaxations

### v3

@@ Abstract
[Abstract] Machine-learning interatomic potentials have revolutionized materials modeling at the atomic scale. Thanks to these, it is now indeed possible to perform simulations of ab initio quality over very large time and length scales. More recently, various universal machine-learning models have been proposed as an out-of-box approach avoiding the need to train and validate specific potentials for each particular material of interest. In this paper, we review and evaluate four different universal machine-learning interatomic potentials (uMLIPs), all based on graph neural network architectures which have demonstrated transferability from one chemical system to another. The evaluation procedure relies on data both from a recent verification study of density-functional-theory implementations and from the Materials Project. Through this comprehensive evaluation, we aim to provide guidance to materials scientists in selecting suitable models for their specific research problems, offer recommendations for model selection and optimization, and stimulate discussion on potential areas for improvement in current machine-learning methodologies in materials science.
@@
Materials simulations at the atomic scale are the backbone of computational materials design and discovery. They rely on the Born Oppenheimer approximation, in which the electrons follow the nuclear motion adiabatically, so that the potential governing the nuclei consists of the electronic energies as a function of the nuclear positions, called "potential energy surface" (PES). Knowledge of the PES allows the identification of stable and metastable atomic configurations from minimum energy search, or the determination of materials properties as thermodynamical averages from molecular dynamics simulations 1. The utility of atom-based materials simulations is thus intimately related to the generation of accurate PESs, which has been possible in the last decades thanks to the advent of density-functional theory (DFT) 2, 3, 4, 5. Nonetheless, this ab initio approach relies on the quantum mechanical solution of the electronic problem whose computational cost scales cubically with system size and can therefore become unaffordable in various significant cases of technological interest such as amorphous solids, interfaces, surfaces, etc. At the other end of the simulation approaches, parametrized approximations of the Born-Oppenheimer PES, known as empirical analytical potentials, or force fields, or "classical” interatomic potentials, have been widely used especially for large-scale materials studies 6. Unfortunately, in particular when complex electron interactions are involved (as in chemical reactions or phase transitions) these approaches cannot usually achieve DFT accuracy, and in addition they have limited applicability and transferability. They cannot therefore be considered as a drop-in replacement for standard ab initio methods. In this context, machine-learning interatomic potentials (MLIPs) have emerged as an in-between solution with computational cost similar to the empirical analytical potentials, but with the promise of achieving an accuracy comparable to DFT hence enabling accurate simulations over very large time and length scales 7. The key difference with respect to the empirical potentials is that the interatomic potentials are now directly obtained via a highly non linear fit of a set of input/target data (in general, at the DFT accuracy), without any a priori assumption on their analytical form 8. In the original formulation, building a MLIP consists of generating a dataset of atomic configurations for the specific material under study, and training (and subsequently validating) the MLIP on these data based on the accurate prediction of some target metrics as, e.g., energies, forces, and stresses 8, 9, 10, 11, 12, 13, 14, 15, 16. This process is highly material-dependent, and usually requires a significant human and computational effort 17. Obtaining an accurate description of the PES without the need for costly DFT computations, but also covering all possible chemical and structural spaces, would be the holy grail of MLIPs. In this framework, graph Neural Networks (GNN) have revealed very effective, in particular in message passing and equivariant 15, 18 NN 19. In general, NN-based methods can be particularly useful for generalizability thanks to their property of “learning locally” that makes the resulting potential less material-dependent 8, 9, 20, 21, 22. Indeed, the first “universal” MLIP (uMLIP), the MEGNet 23 model, exploited a graph network architecture. It was trained on 60000 inorganic crystals in their minimum energy configurations from the Materials Project (MP) 24 database, which covers 89 elements of the periodic table and is based on the Perdew-Burke-Enzerhof (PBE) exchange-correlation functional 25. This model could provide the formation energy as well as a number of other properties. In order to predict forces and stresses, various other models were subsequently developed relying on different datasets. In particular, the M3GNet 26 and CHGNet 27 models relying on equivariant graph neural network architectures were both developed using snapshots from DFT relaxations of the MP structures. The publication introducing CHGNet was also the opportunity for releasing the Materials Project Trajectory (MPtrj) dataset 27, including the DFT calculations for more than 1.5 million atomic configurations of inorganic structures. By including magnetic moments in the training properties, CHGNet aimed at a better description of chemical reactions as charged states influence how atoms connect with others through chemical bonds 27. A re-implementation of M3GNet, MatGL, has been built on the Deep Graph Library and on PyTorch in order to improve usability and scalability 28. Recently, the M3GNet-DIRECT model has been released 29 in the MatGL package 28 with 1.3 million structures and 89 elements, aiming to predict more reliably unseen structures thanks to a new sampling scheme 29. The ALIGNN-FF model 30 was developed to model a diverse set of materials with any combination of 89 elements from the periodic table but relying on a different database of inorganic crystals, JARVIS-DFT 31, which is based on the optB88vdW exchange-correlation functional 32. Recently, the MACE-MP-0 33 relying on the MACE architecture 19 was also proposed. It was trained on the MPtrj dataset and showed outstanding performance on an extraordinary range of examples from quantum-chemistry and materials science 33. Two proprietary models relying on very large databases have also been developed: the GNoME 34 model exploiting the NequIP architecture 15 and the PFP model 35, 36 exploiting the TeaNet architecture 37. On the one hand, GNoME was trained on a database obtained from a complex active learning workflow of the original MP data, resulting in a number of inorganic structures 100 times larger than the MPtrj. On the other hand, the PFP was trained on a large dataset which initially included 107 DFT configurations covering 45 elements 35 and was further extended to cover 72 elements at the moment of the publication 36, with an expected rise to 94 including rare-earth elements and actinides. It is worth mentioning that uMLIPs have also been developed specifically for organic molecules 38, 39, 40, and for metal alloys 41. While this rapid growth of uMLIP models promises an extraordinary impact in the materials science community, nevertheless benchmark efforts are crucially needed to assess and control their performance and effective usefulness 42, 43, 44.
@@
In this paper, we conduct a comprehensive review and evaluation of four different GNN-based uMLIPs: M3GNet-DIRECT 29, CHGNet 27, MACE-MP-0 33, and ALIGNN-FF 30. They have demonstrated the possibility of universal interatomic potentials that may not require retraining for new applications. The evaluation uses three different datasets: for the equation of state, we use the set of theoretical structures employed in Ref. 45, for the phonon calculations, we use the crystalline structures considered in Ref. 46 that have been relaxed with Abinit 47, 48, norm-conserving pseudopotentials 49, 50, and the PBEsol exchange-correlation functional 51 while for all the other tests, we use the VASP-relaxed structures from the Materials Project (MP) 24, 52.
@@
The paper is organized as follows. After a brief overview of the techniques employed in the generation of the different uMLIPs examined in this work (Sec. 2.1), we test and discuss the quality and transferability of the chosen uMLIPs in different types of calculations, and using the three different datasets as explained above. The tests include the calculation of the equation of state (Sec. 2.2), the evaluation of formation energies and optimization of structural parameters (Sec. 2.3), and the calculation of phonon bands (Sec. 2.4). The conclusions are provided in Sec. 3 and the details of the methods employed are given in Sec. 4.
@@ 2 Discussion
2.1 Overview of the uMLIP models
In what follows, we give a brief overview of the basic ingredients underlying
the four GNN-based uMLIPs
examined 27, 29, 30, 33.
Depending on the uMLIPs, they might differ for the specific GNN architecture
employed, for the choice of the training set, for the sampling scheme, or for
the properties included in the training. A comprehensive and detailed
explanation is beyond the scopes of the present work, and we refer the
interested reader to the respective original
papers 27, 29, 30, 33 and
related literature 19, 23, 26, 31, 53, 54, 55, 56.
GNN are reviewed, among others, in
Refs. 18, 57, 58, while
Refs. 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 59 address specifically the construction of MLIP
in materials science.
The pre-trained M3GNet-DIRECT 29 model starts from the Materials 3-body Graph Network (M3GNet) architecture 26, which conjugates the flexibility of typical graph techniques with the physically based many-body features (here 3-body) of traditional interatomic potentials like the Tersoff bond-order potential 60.
The many-body interactions (angles) are aggregated to bonds in standard graph convolution steps to update the bond, atom, and global state (properties) information 26.
The main improvement of M3GNet-DIRECT 29 with respect to the original M3GNet 26 is in the sampling of the materials space. This is done starting from the 128-element vector outputs from the M3GNet 26 model trained on the MP formation energies database 24, to which dimensionality reduction, characteristics-sharing-based clustering, and stratification, are subsequently applied, giving the name of DImensionality-Reduced Encoded Clusters with sTratified (DIRECT) sampling to the approach 29. Static DFT calculations on the so-obtained 1.3 million structures produce the final training set 29. The result is an improved performance with respect to the original M3GNet 26, with a better transferability and prediction of configurations with large energies, forces, and stresses 29.
The Crystal Hamiltonian Graph Neural Network (CHGNet 27) is a model pre-trained on the energies, forces, stresses, and magnetic moments from the MPtrj 27 dataset. The GGA/GGA+U mixing compatibility corrections 61, 62, 63 are applied to the energies 27.
In the CHGNet architecture, the angle information is drawn as a pairwise message passing convolution between bonds (bond graph, where bonds are nodes and edges are angles), on top of the bond information which is drawn as a pairwise convolution between atoms (atom graph, where atoms are nodes and edges are bonds) 27.
To constrain the atom features used to predict energy, forces, and stresses by their charge-state
information, the latter is inferred from the magnetic moments and atomic orbital theory before the last convolution layer 27.
In incorporating the site-specific magnetic moments as the input charge states into the GNN, this model aims to capture the chemical interaction variability across different valence states 27. This should be especially important for transition-metal elements, with variable valence states 64.
The pre-trained MACE-MP-0 model 33 relies on MACE 19, an equivariant message passing neural network potential using messages of an order higher than two body. In combining high body order with message passing, MACE achieves superior accuracy as a result of a trade-off between using more layers and increasing the local body order within a single layer 55.
Using 4-body messages 19, the required number of message passing iterations is reduced to two and the steepness of the learning curve is improved 19.
The high-order equivariant message passing neural network model is combined with the atomic cluster expansion (ACE 13, 65), a framework for deriving an efficient body-ordered symmetric polynomial basis
to represent functions of atomic neighborhoods.
ACE naturally extends to
equivariant features and to include variables beyond geometry, such as charges or magnetic moments 66. MACE-MP-0 is pre-trained on the MPtrj 27 dataset, including a variety of magnetic orders 33. The fitting does not explicitly account for the GGA/GGA+U mixing compatibility corrections 61, 62, 63.
The pre-trained ALIGNN-FF model 30 is the extension to derivative prediction and force field (FF) formalism 30 of the Atomistic LIne Graph Neural Network (ALIGNN) models 56, able to capture two and three body
interactions and to output more than 70 materials properties 30. The ALIGNN 56 models
were implemented using the deep graph library (DGL) 67 which allows efficient line graph construction and neural message passing.
As in the case of CHGNet, the message passing occurs both in an atom graph, where atoms are nodes and interatomic bonds are edges, and in a line graph, where bonds are nodes and bond angles are edges 56.
For the atom graph, nine input node features are assigned to each node, based on the atomic species involved: electronegativity, group
number, covalent radius, valence electrons, first ionization
energy, electron affinity, block and atomic volume.
ALIGNN is a part of the Joint Automated Repository for Various Integrated Simulations (JARVIS) infrastructure 31, a set of databases and tools for materials design. The ALIGNN-FF model was trained on the JARVIS-DFT dataset which contains around 75,000 materials and 4 million energy-force entries,
out of which 307,113 were used in the training 30.
For the sake of simplicity, in what follows we use a shorter nomenclature for three out of the four uMLIPs benchmarked: M3GNet-DIRECT 29 will be called M3GNet, MACE-MP-0 33 will be referred to as MACE, and ALIGNN-FF 30 will be termed ALIGNN.
Some key features of the four uMLIPs are listed in Table 1.
| uMLIP label | Model | Version | Model Size | Data Set | Nelements | Data Size |
| ‘CHGNet’ | CHGNet 27 | 0.3.0 68 | 400K 27 | MP-trj 69 | 89 | 1.58M 27 |
| ‘M3GNet’ | M3GNet-DIRECT 29 | 2021.2.8-DIRECT-PES 70 | 1.1M 43 | MPF 24 + DIRECT 29 | 128 | 1.31M 29 |
| ‘MACE’ | MACE-MP-0 33 | 2023-12-03-mace-128-L1_epoch-199 71 72 | 4.7M 44 | MP-trj 69 | 89 | 1.58M 27 |
| ‘ALIGNN’ | ALIGNN-FF 30 | alignnff_wt10 73 74 | 4.0M 75 | JARVIS-DFT 31 | 89 | 75K 30, 31 |
@@ 2.2 Equation of state and comparison with all-electron results
As a first test, we use the protocol for the equation of state (EOS) detailed in Ref. 45, where a high-quality reference dataset of EOS for 960 cubic crystal structures is generated by employing two all-electron (AE) codes. This dataset includes all elements from Z = 1 (hydrogen) to Z = 96 (curium). For each element X, four mono-elemental cubic crystals (unaries structures) are considered, in the face-centered cubic (FCC), body-centered cubic (BCC), simple cubic (SC), and diamond crystal structure, respectively. Besides, for each element, six binary cubic oxides (oxides structures) are included, with chemical formula X2O, XO, X2O3, XO2, X2O5, and XO3, respectively. While in Ref. 45 this dataset is used to gauge the precision and transferability of nine pseudopotential-based ab initio codes, we use it here to assess the four tested uMLIPs.
@@
In Fig. 1, the relative errors on the uMLIPs predictions for the equilibrium volume , the bulk modulus , and its derivative with respect to the pressure obtained from a fit 45 of the EOS, are compared with the analogous errors from selected ab initio methods employed in Ref. 45. FLEUR 76 and WIEN2K 77 are AE codes while VASP 52 implements the projector augmented-wave method (PAW) method 78 with a planewave basis set. The errors reported in Fig. 1 are calculated with respect to the average of the AE methods considered here, i.e., FLEUR 76 and WIEN2K 77.
@@
Furthermore, as in Ref. 45, the EOS computed with two different computational approaches and ( and ) are compared through two metrics and . The first is a renormalized dimensionless version of the metrics introduced in Ref. 79:
@@
where the index runs over the explicit calculations of for the different methods and is the average of over the considered volume range. The second metrics, dependent directly on the physically measurable quantities , , and , captures the relative deviation for each of these three parameters between the two computational approaches and :
@@
where , , and are appropriately chosen weights (see Ref. 45). In Eqs. (1)-(2), is a uMLIP and is the average of the AE methods considered here, i.e., FLEUR 76 and WIEN2K 77. Heatmaps of the periodic table with the values of the comparison metrics and obtained with all the different uMLIPs are reported in Supplemental Material Figs. S1-S4.
@@
It should be noted that most structures in the dataset used in this EOS test are not stable in nature, so that they were not likely included in the dataset used to train the uMLIPs. As a consequence, it is not surprising that the uMLIPs are not able to predict the correct energy versus volume curve for a significant fraction of the systems (Figs. S1-S4). Nevertheless, even for those systems for which a physical EOS is obtained, the precision and transferability is still far from the one that can be achieved with state-of-the-art pseudopotential-based ab initio techniques, as can be observed in Fig. 1. This is a very stringent test for uMLIPs, especially given the effort made in Ref. 45 to improve the precision and the transferability of pre-existent pseudopotential tables used for ab initio calculations. Yet these results suggest that uMLIPs predictions should be taken with some caution and, if possible, validated a posteriori via ab initio calculations, especially if the chemical/physical environment under study is not properly included in the training dataset. The precision of uMLIPs can be improved after retraining the model by including additional ab initio data capturing the chemical/physical configurations under investigation 43, 44. However, despite their relevance, these topics are beyond the scope of the present work and are left for future investigation.
2.3 Structural optimization and formation energies